# UVE: Are MLLMs Unified Evaluators for AI-Generated Videos?

### v2

@@ Abstract
With the rapid growth of video generative models (VGMs), it is essential to develop reliable and comprehensive automatic metrics for AI-generated videos (AIGVs). Existing methods either use off-the-shelf models optimized for other tasks or rely on human assessment data to train specialized evaluators. These approaches are constrained to specific evaluation aspects and are difficult to scale with the increasing demands for finer-grained and more comprehensive evaluations. To address this issue, this work investigates the feasibility of using multimodal large language models (MLLMs) as a unified evaluator for AIGVs, leveraging their strong visual perception and language understanding capabilities. To evaluate the performance of automatic metrics in unified AIGV evaluation, we introduce a benchmark called UVE-Bench. UVE-Bench collects videos generated by state-of-the-art VGMs and provides pairwise human preference annotations across 15 evaluation aspects. Using UVE-Bench, we extensively evaluate 16 MLLMs. Our empirical results suggest that while advanced MLLMs (e.g., Qwen2VL-72B and InternVL2.5-78B) still lag behind human evaluators, they demonstrate promising ability in unified AIGV evaluation, significantly surpassing existing specialized evaluation methods. Additionally, we conduct an in-depth analysis of key design choices that impact the performance of MLLM-driven evaluators, offering valuable insights for future research on AIGV evaluation. The code is available at https://github.com/bytedance/UVE.
Machine Learning, ICML
@@ 1 Introduction
Video generative models (VGMs) have rapidly evolved in recent years, greatly aiding visual content creation in numerous fields. However, even current state-of-the-art (SOTA) VGMs (Brooks et al., 2024; Polyak et al., 2024; Kong et al., 2024; Kuaishou., 2024) still suffer from issues like incorrect subject structure, unnatural motion, and imperfect alignment with the text prompt. Consequently, it is crucial to develop automatic metrics that can effectively identify these imperfections in modern AI-generated videos (AIGVs), so as to facilitate the advancement of VGMs and enhance the application of AIGVs in real-world scenarios.
@@
Existing AIGV evaluation metrics fall into two categories. The first employs off-the-shelf models (e.g., CLIP (Radford et al., 2021) and DINO (Caron et al., 2021)) and designs heuristic rules to assess various specific aspects of AIGVs (Huang et al., 2024; Liu et al., 2024b). The second category involves collecting human assessments for specific aspects of interests and training models to imitate these assessments (He et al., 2024; Bansal et al., 2024; Zhang et al., 2024d; Wu et al., 2024a; Wang et al., 2024a). Such specifically trained automatic evaluators have shown better correlation with human evaluations compared to the off-the-shelf models. However, with the rapid development of VGMs and AIGV applications, there is an urgent need for finer-grained and more comprehensive evaluations. Continuously collecting human assessments and training models to accommodate emerging evaluation aspects and frequently changing standards is cost-intensive and difficult to scale.
@@
Unlike existing automatic metrics, we humans can assess any aspect of AIGVs as long as a proper guideline is provided. This ability stems from our robust visual perception and language understanding. Leveraging the knowledge learned from vast amounts of visual and language data, the latest multimodal large language models (MLLMs) (Wang et al., 2024c; OpenAI, 2023; Chen et al., 2024b) also exhibit strong ability in joint vision-language understanding. This progress naturally raises a question: “Can MLLMs be utilized as a unified AIGV evaluator like humans?
@@
To address this question, we propose an approach to unify AIGV evaluations by prompting pre-trained MLLMs and mapping their outputs to evaluation results. This method enables zero-shot evaluation of any aspect of AIGV by simply modifying the prompts. It supports both single video ratings and video pair comparisons. While using MLLMs for AIGV evaluation is a straightforward concept and has been explored in previous works (He et al., 2024; Bansal et al., 2024; Wang et al., 2024a), these studies are limited by evaluating only a few aspects and relying on human-annotated ratings to train the MLLMs.
@@
To evaluate the capability of MLLMs as unified AIGV evaluators, we require a benchmark that (1) encompasses a broad range of AIGV aspects, (2) provides accurate human evaluations as references, and (3) includes AIGVs that highlight the weaknesses of state-of-the-art VGMs. Since no existing AIGV dataset meets all these criteria, we introduce a new benchmark called UVE-Bench (short for Unified Video Evaluation). UVE-Bench has three key features: First, it covers a wide range of 15 evaluation aspects. Second, it provides human annotation in the form of pairwise video preference, avoiding the inconsistent standards of absolute ratings between humans (Wu et al., 2024b), and can be used to evaluate both single video ratings and video pair comparisons. Third, the videos in UVE-Bench are generated by the latest VGMs (e.g., MovieGenVideo (Polyak et al., 2024) and HunyuanVideo (Kong et al., 2024)), challenging the evaluators to identify the weaknesses of such videos.
@@
Based on UVE-Bench, we conduct an extensive evaluation of 16 MLLMs. Our results indicate that unified evaluator powered by advanced MLLMs, such as Qwen2-VL-72B (Wang et al., 2024c) and InternVL2.5-78B (Chen et al., 2024b), indeed show promising abilities in evaluating various AIGV aspects, outperforming existing approaches that focus on specific aspects. However, there remains a significant gap between these MLLMs and human evaluators, particularly in aspects requiring a fine-grained understanding of temporal dynamics in videos. Additionally, we performed a series of analytical studies on the design choices that impact the performance of our MLLM-driven AIGV evaluation framework, providing valuable insights for future research in the field.
The contributions of this work are summarized as follows:
  1. 1.
    We introduce a unified approach to evaluate any aspect of AIGV using pre-trained MLLMs.
  2. 2.
    We propose UVE-Bench, a comprehensive benchmark to assess the capability of unified AIGV evaluation.
  3. 3.
@@
    We conduct in-depth analysis on the pros and cons of MLLMs in unified AIGV evaluation and the key design choices that impact their performance.
2 Related Work
2.1 Video Generative Models
The emergence of diffusion models (Ho et al., 2020; Song et al., 2021) has revolutionized generative applications, spanning from image synthesis (Saharia et al., 2022; Ramesh et al., 2022; Rombach et al., 2022) to video generation (Gupta et al., 2023; Yu et al., 2023; Ho et al., 2022a, b). On the one hand, recent VGMs like SORA (Brooks et al., 2024), Kling (Kuaishou., 2024), and MovieGen (Polyak et al., 2024) have achieved remarkable success in producing highly realistic videos that closely resemble real ones. On the other hand, the condition signal for video generation has been expanded from text (T2V) to RGB images (I2V) (Li et al., 2023c; Chen et al., 2023a) and other videos (V2V) (Zhao et al., 2024). This work specifically focuses on evaluating text-conditioned video generation.
2.2 Multimodal Large Language Models
Multimodal large language models (OpenAI, 2024; Wang et al., 2024c; Li et al., 2024, 2023a) integrate LLMs with additional perception modules, enabling them to comprehend information beyond the text modality. With advancements in LLM backbones and the increasing scale of multimodal training, MLLMs are showing promising performance in understanding real-world images (Liu et al., 2023a, 2024a; Gao et al., 2024; Wang et al., 2024b) and videos (Zhang et al., 2024b; Li et al., 2023b; Lin et al., 2023). However, understanding AI-generated visual content presents unique challenges compared to real images, and the application of MLLMs in this field is still in its early stages.
2.3 AI-Generated Video Evaluation
Despite the progress in VGMs, they still struggle to consistently generate high-quality videos, which necessitate a comprehensive and fine-grained evaluation. Initial evaluation approaches (Huang et al., 2024; Liu et al., 2024b) employ off-the-shelf models to evaluate specific aspects, such as CLIP (Radford et al., 2021) for video-text alignment and DINO (Caron et al., 2021) for frame-to-frame consistency. With the advancement of MLLMs, recent studies have incorporated these models in AIGV evaluation (He et al., 2024; Wu et al., 2024a; Wang et al., 2024a; Wu et al., 2024c; Bansal et al., 2024). However, they are constrained to a fixed set of evaluation aspects and typically rely on human annotations to further fine-tune the MLLMs for AIGV evaluation. By contrast, in this work explores using MLLMs as unified evaluators for any AIGV aspect by leveraging their inherent vision-language understanding capabilities, eliminating the need for human annotations.
Single Video Rating <video> Watch the above frames of an AI-generated video and evaluate <aspect-specific description> Complete your evaluation by answering this question: <aspect-specific question>? <answer prompt> Video Pair Comparison
The first video: <video>
The second video: <video> Watch the above two AI-generated videos and evaluate <aspect-specific description> Complete your evaluation by answering this question: Which video is <aspect-specific question>? You should make your judgment based on the following rules: <instructions on how to make the choice> Now give your judgment:
Table 1: MLLM prompting templates for unified AIGV evaluation.
@@ 3 Unifying AIGV Evaluation with MLLMs
In this section, we introduce how to utilize a single MLLM to evaluate any aspect of AIGVs, focusing on both single video rating and video pair comparison.
@@ 3.2 Method
Research on image quality evaluation (Zhang et al., 2024e; Lin et al., 2024) has demonstrated that, given an appropriate prompt, the generative likelihood of certain tokens (e..g, yes/no or good/poor) by MLLMs can be used an effective indicator of image visual quality and image-text alignment. Motivated by this, we formulate single video rating in a unified manner as:
@@
A straightforward way to perform video pair comparison is directly feeding the video pair into MLLMs, as the latest MLLMs (Wang et al., 2024c; Li et al., 2024; Chen et al., 2024b) already support interleaved vision-language input with multiple videos. Specifically, we feed the video pair, along with the corresponding T2V prompts if needed, into the MLLM and prompts it to make a choice from $\mathbf{O}$:
@@
Table 1 illustrates the prompting templates for single video rating and video pair comparison, which can be adapted to any evaluation aspect by modifying the aspect-specific prompt. The specific prompts being used for different aspects are provided in the supplementary material.
Figure 1: Overview of UVE-Bench. (a) The distribution of video sources. (b) The distribution of data example over 15 fine-grained AIGV evaluation aspects. (c) The distribution of human preference over the four categories. (d) Data examples illustrating how to evaluate both single video rating and video pair comparison using the human preference annotations. More examples can be found in Appendix B.4.
@@ 4 UVE-Bench
As illustrated in Figure 1, UVE-Bench is designed to assess AIGV automatic evaluation methods. This section details the process of video collection (§4.1), the evaluation aspects covered by UVE-Bench (§4.2), the criterion for assessing the performance of automatic AIGV evaluators using human preferences (§4.3), the data annotation procedure (§4.4) and a comparison with existing AIGV datasets (§4.5).
Video Source Prompt Source Resolution Duration FPS Release Date Open Source Gen-3  (Runway., 2024) VideoGen-Eval 1280$\times$768 5.3 24 2024.06 Kling 1.5  (Kuaishou., 2024) VideoGen-Eval 1920$\times$1080 5.1 30 2024.09 LUMA 1.6  (LumaLabs., 2024) VideoGen-Eval 1360$\times$752 5.0 24 2024.09 Movie Gen Video  (Polyak et al., 2024) Movie Gen 848$\times$480 5.0 24 2024.10 Open-Sora 1.2  (Zheng et al., 2024) VideoGen-Eval Movie Gen ShutterStock 1280$\times$720 848$\times$480 4.3 5.3 24 2024.06 CogVideoX-5B (Yang et al., 2024b) VideoGen-Eval Movie Gen ShutterStock 720$\times$480 6.1 8 2024.08 HunyuanVideo (Kong et al., 2024) Movie Gen ShutterStock 960$\times$544 5.4 24 2024.12 Mochi 1 (Team, 2024) Movie Gen ShutterStock 848$\times$480 5.4 30 2024.10 ShutterStock ShutterStock 600$\times$316 898$\times$506 596$\times$336 5.0 24, 25 30, 60 - -
@@ 4.1 Video Collection
The videos in UVE-Bench are created using eight of the latest video generative models via text-to-video generation. To ensure diverse video content, we select text prompts from three sources: (1) VideoGen-Eval (Zeng et al., 2024) provides over 700 prompts, targeting various application scenarios (e.g., animation and advertisement) and key capabilities (e.g., text alignment and motion diversity) of video generation. (2) Movie Gen Video Bench (Polyak et al., 2024) offers 1,003 prompts covering various categories of video content (e.g., human activity, animals, and scenery). (3) We also modify a small portion of video captions from the ShutterStock [^1] https://www.shutterstock.com/ platform to generate videos with different degrees of Light & Color change (an evaluation aspect that will be discussed in §4.2).
For VideoGen-Eval, we directly collect videos generated by Gen-3, Kling 1.5, LUMA 1.6, Open-Sora 1.2, and CogVideoX-5B, as released by the authors of Zeng et al. (2024). For Movie Gen Video Bench and ShutterStock, we generate videos using Open-Sora 1.2, CogVideoX-5B, Mochi 1 and HunyuanVideo. For the Movie Gen Video model, we directly collect the generated videos released by Polyak et al. (2024) and select the first five seconds of the videos. Additionally, we collect some non-AIGV videos related to Light & Color from ShutterStock. As shown in Table 2, the videos have a duration of around 5.0 seconds, with resolution ranging from $596\times 336$ (360p level) to $1280\times 720$ (720p level) and $1920\times 1080$ (1080p level).
@@ 4.2 Evaluation Aspects
Inspired by previous work in AIGV evaluation (Huang et al., 2024; Liu et al., 2023b, 2024b), we identify four fundamental evaluation aspects to build our benchmark: Static Quality, Temporal Quality, Dynamic Degree and Video-Text Alignment, each of them is further divided into several subaspects.
@@ 4.3 Evaluation Criterion
As shown in Figure 1, each example in UVE-Bench consists of three elements: a pair of videos $\{\mathcal{V}_{1},\mathcal{V}_{2}\}$, the evaluation aspect $a$ and corresponding human preference annotation $\mathcal{P}\in\mathbf{O}=\{``\mathcal{V}_{1}\text{better}",``\mathcal{V}_{2}\text{better}",``\text{same good}",``\text{same bad}"\}$.
@@
$\mathcal{A}^{\text{single}}=\begin{cases}\mathbf{1}(\mathcal{S}_{1}>\mathcal{S}_{2})&\text{if }\mathcal{P}=``\mathcal{V}_{1}\text{better}"\\ \mathbf{1}(\mathcal{S}_{1}<\mathcal{S}_{2})&\text{elif }\mathcal{P}=``\mathcal{V}_{2}\text{better}"\\ f_{c}(\mathcal{S}_{1}|\beta)\cdot f_{c}(\mathcal{S}_{2}|\beta)&\text{elif }\mathcal{P}=``\text{same good}"\\ f^{{}^{\prime}}_{c}(\mathcal{S}_{1}|\alpha)\cdot f_{c}(\mathcal{S}_{2}|\alpha)&\text{elif }\mathcal{P}=``\text{same bad}"\\ \end{cases}$
Here, $\mathbf{1}(\cdot)\in\{0,1\}$ is a binary indicator function and $\alpha<\beta$ are the thresholds for “bad” and “good”, respectively. $f_{c}(\mathcal{S}|\beta)$ is a piecewise function that equals to 1 when $\mathcal{S}\in[\beta,1]$, indicating that the model evaluation aligns with the human judgment of “same good”. When $\mathcal{S}\in[0,\beta)$, $f_{c}(\mathcal{S}|\beta)$ exponentially decays from 1 to 0, indicating a gradual deviation from “same good”. Similarly, $f^{{}^{\prime}}_{c}(\mathcal{S}|\alpha)$ is constructed in a reversed manner to assess whether $\mathcal{S}$ agrees with the human judgment of “same bad”. Details of $f_{c}(\mathcal{S}|\alpha)$ and $f^{{}^{\prime}}_{c}(\mathcal{S}|\beta)$ can be found in Appendix A.2.

### v4

@@ Abstract
With the rapid growth of video generative models (VGMs), it is essential to develop reliable and comprehensive automatic metrics for AI-generated videos (AIGVs). Existing methods either use off-the-shelf models optimized for other tasks or rely on human assessment data to train specialized evaluators. These approaches are constrained to specific evaluation aspects and are difficult to scale with the increasing demands for finer-grained and more comprehensive evaluations. To address this issue, this work investigates the feasibility of using multimodal large language models (MLLMs) as a unified evaluator for AIGVs, leveraging their strong visual perception and language understanding capabilities. To evaluate the performance of automatic metrics in unified AIGV evaluation, we introduce a benchmark called UVE-Bench. UVE-Bench collects videos generated by state-of-the-art VGMs and provides pairwise human preference annotations across 15 evaluation aspects. Using UVE-Bench, we extensively evaluate 18 MLLMs. Our empirical results suggest that while advanced MLLMs (e.g., Qwen2VL-72B and InternVL2.5-78B) still lag behind human evaluators, they demonstrate promising ability in unified AIGV evaluation, significantly surpassing existing specialized evaluation methods. Additionally, we conduct an in-depth analysis of key design choices that impact the performance of MLLM-driven evaluators, offering valuable insights for future research on AIGV evaluation. [Uncaptioned image] Code: https://github.com/bytedance/UVE, [Uncaptioned image] Data: https://huggingface.co/datasets/lyx97/UVE-Bench.
@@ 1 Introduction
Video generative models (VGMs) have rapidly evolved in recent years, greatly aiding visual content creation in numerous fields. However, even current state-of-the-art (SOTA) VGMs [3, 46, 25, 27] still suffer from issues like incorrect subject structure, unnatural motion, and imperfect alignment with the text prompt. Consequently, it is crucial to develop automatic metrics that can effectively identify these imperfections in modern AI-generated videos (AIGVs), so as to facilitate the advancement of VGMs and enhance the application of AIGVs in real-world scenarios.
@@
Existing AIGV evaluation metrics fall into two categories. The first employs off-the-shelf models (e.g., CLIP [47] and DINO [5]) and designs heuristic rules to assess various specific aspects of AIGVs [21, 40]. The second category involves collecting human assessments for specific aspects of interests and training models to imitate these assessments [17, 2, 75, 61, 56]. Such specifically trained automatic evaluators have shown better correlation with human evaluations compared to the off-the-shelf models. However, with the rapid development of VGMs and AIGV applications, there is an urgent need for finer-grained and more comprehensive evaluations. Continuously collecting human assessments and training models to accommodate emerging evaluation aspects and frequently changing standards is cost-intensive and difficult to scale.
@@
Unlike existing automatic metrics, we humans can assess any aspect of AIGVs as long as a proper guideline is provided. This ability stems from our robust visual perception and language understanding. Leveraging the knowledge learned from vast amounts of visual and language data, the latest multimodal large language models (MLLMs) [58, 44, 9] also exhibit strong ability in joint vision-language understanding. This progress naturally raises a question: “Can MLLMs be utilized as a unified AIGV evaluator like humans?
@@
To address this question, we propose an approach to unify AIGV evaluations by prompting pre-trained MLLMs and mapping their outputs to evaluation results (§2). This method enables zero-shot evaluation of any aspect of AIGV by simply modifying the prompts. It supports both single video ratings and video pair comparisons. While using MLLMs for AIGV evaluation is a straightforward concept and has been explored in previous works [17, 2, 56], these studies are limited by evaluating only a few aspects and relying on human-annotated ratings to train the MLLMs.
@@
To evaluate the capability of MLLMs as unified AIGV evaluators, we require a benchmark that (1) encompasses a broad range of AIGV aspects, (2) provides accurate human evaluations as references, and (3) includes AIGVs that highlight the weaknesses of state-of-the-art VGMs. Since no existing AIGV dataset meets all these criteria, we introduce a new benchmark called UVE-Bench (short for Unified Video Evaluation; see §3). UVE-Bench has three key features: First, it covers a wide range of 15 evaluation aspects. Second, it provides human annotation in the form of pairwise video preference, avoiding the inconsistent standards of absolute ratings between humans [62], and can be used to evaluate both single video ratings and video pair comparisons. Third, the videos in UVE-Bench are generated by the latest VGMs (e.g., MovieGenVideo [46] and HunyuanVideo [25]), challenging the evaluators to identify the weaknesses of such videos.
@@
Based on UVE-Bench, we conduct an extensive evaluation of 18 MLLMs. Our results indicate that unified evaluator powered by advanced MLLMs, such as Qwen2-VL-72B [58] and InternVL2.5-78B [9], indeed show promising abilities in evaluating various AIGV aspects, outperforming existing approaches that focus on specific aspects. However, there remains a significant gap between these MLLMs and human evaluators, particularly in aspects requiring a fine-grained understanding of temporal dynamics in videos. Additionally, we perform a series of analytical studies on the design choices that impact the performance of our MLLM-driven AIGV evaluation framework, providing valuable insights for future research in the field.
@@
The contributions of this work are: (1) We introduce a unified approach to evaluate any aspect of AIGV using pre-trained MLLMs. (2) We propose UVE-Bench, a comprehensive benchmark to assess the capability of unified AIGV evaluation. (3) We conduct in-depth analysis on the pros and cons of MLLMs in unified AIGV evaluation and the key design choices that impact their performance.
@@ 2 Unifying AIGV Evaluation with MLLMs
As shown in Fig. 1, the unified evaluator is designed to handle all aspects of AIGV evaluation, eliminating the need for specialized models tailored to individual aspects. In this section, we describe how this framework can be applied to both single video rating and video pair comparison.
@@ 2.1 Problem Formulation
Figure 1: Illustration of MLLM-based unified evaluator and specialized evaluators.
Table 1: MLLM prompting templates for unified AIGV evaluation.
Single Video Rating <video> Watch the above frames of an AI-generated video and evaluate <aspect-specific description>
Complete your evaluation by answering this question: <aspect-specific question>? <answer prompt> Video Pair Comparison
The first video: <video>
The second video: <video> Watch the above two AI-generated videos and evaluate <aspect-specific description>
Complete your evaluation by answering this question: Which video is <aspect-specific question>? You should make your judgment based on the following rules: <instructions on how to make the choice>
Now give your judgment:
@@ 2.2 Method
Research on image quality evaluation [76, 35] has demonstrated that, given an appropriate prompt, the generative likelihood of certain tokens (e..g, yes/no or good/poor) by MLLMs can be used an effective indicator of image visual quality and image-text alignment. Motivated by this, we formulate single video rating in a unified manner as:
@@
A straightforward way to perform video pair comparison is directly feeding the video pair into MLLMs, as the latest MLLMs [58, 29, 9] already support interleaved vision-language input with multiple videos. Specifically, we feed the video pair, along with the corresponding T2V prompts if needed, into the MLLM and prompts it to make a choice from $\mathbf{O}$:
@@
Tab. 2.1 illustrates the prompting templates for single video rating and video pair comparison, which can be adapted to any evaluation aspect by modifying the aspect-specific prompt. The specific prompts being used for different aspects are provided in the supplementary material.
Figure 2: Overview of UVE-Bench. (a) The distribution of video sources. (b) The distribution of data example over 15 fine-grained AIGV evaluation aspects. (c) The distribution of human preference over the four categories. (d) Data examples illustrating how to evaluate both single video rating and video pair comparison using the human preference annotations. More examples can be found in Appendix D.6.
@@ 3 UVE-Bench
As illustrated in Fig. 2, UVE-Bench is designed to assess AIGV automatic evaluation methods. This section details the process of video collection (§3.1), the evaluation aspects covered by UVE-Bench (§3.2), the criterion for assessing the performance of automatic AIGV evaluators using human preferences (§3.3), the data annotation procedure (§3.4) and a comparison with existing AIGV datasets (§3.5).
@@ 3.1 Video Collection
The videos in UVE-Bench are created using eight of the latest video generative models via text-to-video generation. To ensure diverse video content, we select text prompts from three sources: (1) VideoGen-Eval [70] provides over 700 prompts, targeting various application scenarios (e.g., animation and advertisement) and key capabilities (e.g., text alignment and motion diversity) of video generation. (2) Movie Gen Video Bench [46] offers 1,003 prompts covering various categories of video content (e.g., human activity, animals, and scenery). (3) We also modify a small portion of video captions from the ShutterStock [^1] https://www.shutterstock.com/ platform to generate videos with different degrees of Light & Color change (an evaluation aspect that will be discussed in §3.2).
For VideoGen-Eval, we directly collect videos generated by Gen-3, Kling 1.5, LUMA 1.6, Open-Sora 1.2, and CogVideoX-5B, as released by the authors of Zeng et al. [70]. For Movie Gen Video Bench and ShutterStock, we generate videos using Open-Sora 1.2, CogVideoX-5B, Mochi 1 and HunyuanVideo. For the Movie Gen Video model, we directly collect the generated videos released by Polyak et al. [46] and select the first five seconds of the videos. Additionally, we collect some non-AIGV videos related to Light & Color from ShutterStock. As shown in Tab. 2, the videos have a duration of around 5.0 seconds, with resolution ranging from $596\times 336$ (360p level) to $1280\times 720$ (720p level) and $1920\times 1080$ (1080p level).
@@ 3.2 Evaluation Aspects
Inspired by previous work in AIGV evaluation [21, 39, 40], we identify four fundamental evaluation aspects to build our benchmark: Static Quality, Temporal Quality, Dynamic Degree and Video-Text Alignment, each of them is further divided into several subaspects.
@@ 3.3 Evaluation Criterion
As shown in Fig. 2, each example in UVE-Bench consists of three elements: a pair of videos $\{\mathcal{V}_{1},\mathcal{V}_{2}\}$, the evaluation aspect $a$ and corresponding human preference annotation $\mathcal{P}\in\mathbf{O}=\{``\mathcal{V}_{1}\text{better}",``\mathcal{V}_{2}\text{better}",``\text{same good}",``\text{same bad}$"\}.
@@
[…]
Here, $\mathbf{1}(\cdot)\in\{0,1\}$ is a binary indicator function and $\alpha<\beta$ are the thresholds for “bad” and “good”, respectively. $f_{c}(\mathcal{S}|\beta)$ is a piecewise function that equals to 1 when $\mathcal{S}\in[\beta,1]$, indicating that the model evaluation aligns with the human judgment of “same good”. When $\mathcal{S}\in[0,\beta)$, $f_{c}(\mathcal{S}|\beta)$ exponentially decays from 1 to 0, indicating a gradual deviation from “same good”. Similarly, $f^{{}^{\prime}}_{c}(\mathcal{S}|\alpha)$ is constructed in a reversed manner to assess whether $\mathcal{S}$ agrees with the human judgment of “same bad”. Details of $f_{c}(\mathcal{S}|\alpha)$ and $f^{{}^{\prime}}_{c}(\mathcal{S}|\beta)$ can be found in Appendix C.2.